OpenStudy (anonymous):

Prove: sin theta - sin theta * cos^2 theta = sin^3 theta. help really don't get this!!! please!!

OpenStudy (anonymous):

anyone......

OpenStudy (anonymous):

@ganeshie8 @mathmale

OpenStudy (shamim):

Sin theta(1-cos^theta) =sin theta*sin^2theta =?

ganeshie8 (ganeshie8):

start with left hand side : \[\sin\theta - \sin\theta *\cos^2\theta\]

ganeshie8 (ganeshie8):

factor out \(\sin\theta\) from both the terms : \[\sin\theta - \sin\theta *\cos^2\theta = \sin\theta[1- \cos^2\theta ]\]

OpenStudy (anonymous):

okay got it

ganeshie8 (ganeshie8):

good, use the identity \(\sin^2\theta + \cos^2\theta = 1\) and write \(1-\cos^2\theta\) as \(\sin^2\theta\)

OpenStudy (anonymous):

how do I do that??

ganeshie8 (ganeshie8):

\[\sin\theta - \sin\theta *\cos^2\theta = \sin\theta[1- \cos^2\theta ] = \sin\theta[\sin^2\theta] = \sin^3\theta\]